OpenStudy (anonymous):

Please help me! plz plz plz Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (lgbasallote):

where are you confused?

OpenStudy (anonymous):

Part 1: y=3x-1. Part 2: It does not matter which point you choose to in the final answer as either point will result in the same answer. I chose the point (2,5) because it looked simpler at first glance.

OpenStudy (lgbasallote):

@Sanguine.Brother ..???

OpenStudy (anonymous):

Well, you are given two points. You can use those two points to find the slope of the line that passes between them using the equation (5-(-4))/(2-(-1)) which should give you 3. From there we know the equation will be y=3x+b and you just pick one of the points to plug in y and x in order to find b in the general equation.

OpenStudy (lgbasallote):

are these questions related??

OpenStudy (lgbasallote):

oh wait..it'syour answer lol

OpenStudy (lgbasallote):

how did you get y = 3x - 1?

OpenStudy (anonymous):

yeap am a little confused ;//

OpenStudy (lgbasallote):

lol it's correct though..just wanted to know which formula :)

OpenStudy (anonymous):

Sorry, let me try to explain a little more. So, to find the slope of the line that passes between those two points you need to subtract the first y-coordinate from the second y-coordinate (5-(-4)) and divide the answer you get (9) by the quotient resulting from subtracting the first x-coordinate from the second x-coordinate (2-(-1) which would be 3. So 9/3=3 (the slope of the line). With point-slope form we know that y=3x+b, with 3 being the slope we found and b being the y-intercept we need to find. To do this, we pick one of the points, either (-1,-4) or (2,5) and plug them into the equation and solve for b, which in this case will be -1.

OpenStudy (anonymous):

this where I got stuck y-5=3x-2 y=3x-2+5 meabe I am using the wrong formula??

OpenStudy (anonymous):

You are not necessarily using the wrong formula, but you are using a different one incorrectly. I used the slope-intercept form to get my answer while you are using point-slope form. To use point-slope form, the equation would be y-5=3(x-2). The slope is multiplying the difference between the x-coordinates.

OpenStudy (lgbasallote):

the difference is that 3 gets distributed 3(x-2) = 3(x) + 3(-2)

OpenStudy (anonymous):

so am good?? I distributed 3 and then y-5=3x-2 y=3x-2+5 y=3x-7 that my final answer is it right?

OpenStudy (lgbasallote):

you did not distribute 3...look at what i said 3(x+2) = 3(x) + 3(-2) what is 3(x)?

OpenStudy (anonymous):

;;;(( now am confused! sorry ;(( don't kill me

OpenStudy (lgbasallote):

confused where?

OpenStudy (anonymous):

ok y-5=3(x-2) I distributed there so now I have y-5=3x-2 am I correct??

OpenStudy (lgbasallote):

no... 3(x+2) is NOT equal to 3x - 2

OpenStudy (lgbasallote):

just follow what im saying.. 3(x-2) = 3(x) + 3(-2) what's 3(x)?

OpenStudy (anonymous):

I have no idea o.O??

OpenStudy (lgbasallote):

that is read as 3 times x what is 3 times x?

OpenStudy (anonymous):

O.o??

OpenStudy (lgbasallote):

2 times x is 2x 4 times x is 4x so what's 3 times x?

OpenStudy (anonymous):

3x

OpenStudy (lgbasallote):

correct..

OpenStudy (lgbasallote):

3(x-2) = 3(x) + 3(-2) 3(x) = 3x 3(-2) = ?

OpenStudy (lgbasallote):

what is 3(-2)?

OpenStudy (anonymous):

-6

OpenStudy (lgbasallote):

so 3(x-2) is....?

OpenStudy (anonymous):

3x-(-6) ??

OpenStudy (lgbasallote):

no..it's 3x + (-6) remember i wrote 3(x-2) = 3x + 3(-2)

OpenStudy (lgbasallote):

simplify 3x + (-6)..

OpenStudy (anonymous):

thats where I got stuck

OpenStudy (anonymous):

isnt it y-5=3x(-6) ?

OpenStudy (anonymous):

I am super stuck in the deep of the bottom of the ocean!

OpenStudy (lgbasallote):

2a + (-1) = 2a - 1 4b + (-2) = 4b - 2 so what 3x + (-6)?

OpenStudy (anonymous):

3x-6 ?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

so now what is the equation?

OpenStudy (anonymous):

y-5=3x-6 ?

OpenStudy (anonymous):

now I have to simplify that?

OpenStudy (lgbasallote):

yes..add 5 to both sides

OpenStudy (anonymous):

y-5=3x-11

OpenStudy (anonymous):

??

OpenStudy (lgbasallote):

how did it become 11 o.O it was 6 just a minute ago :O

OpenStudy (anonymous):

no wait is y=3x-11

OpenStudy (anonymous):

oh umm -6 +5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

o.O??

OpenStudy (anonymous):

O.o??

OpenStudy (lgbasallote):

will you first right the equation? the one with the 6? dont jumpsteps...

OpenStudy (anonymous):

y-5=3x-6 and I add 5 both sides y=3x-11

OpenStudy (lgbasallote):

-6 + 5 is NOT -11

OpenStudy (anonymous):

the six (-6)

OpenStudy (lgbasallote):

try again... -6 + 5 is like 5 - 6

OpenStudy (anonymous):

-1 o.O?

OpenStudy (anonymous):

my bad o.o

OpenStudy (lgbasallote):

-1 is correct.

OpenStudy (lgbasallote):

y = 3x - 1 just like what sanguine said

OpenStudy (anonymous):

yeap, so my final answer is......ta ta tan! y=3x-1

OpenStudy (anonymous):

??

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

that's for part 1

OpenStudy (anonymous):

the very final answer??

OpenStudy (anonymous):

just kidding thank you so much!!!!

OpenStudy (anonymous):

ur very patient. and kind Thank u for ur help

OpenStudy (lgbasallote):

\[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

3 billion!

OpenStudy (anonymous):

jk!